OpenStudy (anonymous):

A circle has a radius of 7. find the diameter and area a. d= 14 a= 49 pie b. d=7 a=49 pie c. d=14 a= 7pie

OpenStudy (anonymous):

d=2*r area =pi*r^2

OpenStudy (anonymous):

so a?

OpenStudy (anonymous):

yes